# Hybrid thermalization in the large $N$ limit

### v1

@@ 4 Global equilibrium maximizes the total entropy for the isolated hybrid system
We demonstrate this for an isolated system $\mathcal{F}$ with two subsectors $\mathcal{U}$ and $\tilde{\mathcal{U}}$ defined by arbitrary equations of states, $P(\varepsilon)$ and $\tilde{P}(\tilde{\varepsilon})$, see Fig. 1. The physical background metric for the full system is the Minkowski metric
@@
×$\qquad\times a(\lambda)b(\lambda)^{d-1}\tilde{a}(\lambda)\tilde{b}(\lambda)^{d-1}.$
is a constant [^2] Since we are considering homogeneous configurations, the total energy density should be held fixed and the total entropy density should be maximized. It will be clear from the proof that it is sufficient to prove the maximization of entropy in global thermal equilibrium in homogeneous equilibrium as the proof can be implemented locally.. We recall that $a(\lambda)$, $b(\lambda)$, $\tilde{a}(\lambda)$ and $\tilde{b}(\lambda)$ should be obtained by solving the coupling equations sec. 3.2 (with $\phi=0$) which are fully specified by the subsystem temperatures, $T_{1}(\lambda)$ and $T_{2}(\lambda)$.
@@
For illustration, we can consider the case of two subsystems with conformal equations of states $\epsilon=n_{1}T_{1}^{3}$ and $\tilde{\epsilon}=n_{2}T_{2}^{3}$ in $2+1-$dimensions. We choose $n_{1}=10$ and $n_{2}=1$. In the left panel in Fig. 2, we note that while one of the subsystem entropies monotonically increases with its effective temperature, that of the other decreases as a consequence of the fixed total energy. With the total energy $\mathcal{E}$ kept constant and setting $\lambda=T_{1}$, the total entropy $\mathcal{S}$ has been plotted alongside $T_{a}-T_{b}$ as a function of $T_{1}$ in the right panel of Fig. 2. We observe that $T_{a}-T_{b}$ indeed vanishes exactly where $\mathcal{S}$ is maximum. We also note that the total entropy is a concave function and has only one maximum corresponding to the global equilibrium.
@@ 5.1 Hybrid BRSSS setup
[…]
@@
where $\tau_{\pi}(\tilde{\tau}_{\pi})$ and $\eta(\tilde{\eta})$ are the relaxation time and shear viscosity, respectively. The shear tensors, $\sigma^{\mu\nu}$ and $\tilde{\sigma}^{\mu\nu}$, are defined as
@@
In the homogeneous thermal relaxation process in $2+1$-dimensions, the full system is governed by fourteen equations; namely, one Ward identity for each subsystem (the conservation of the energy-momentum tensor gives only one non-vanishing equation for each subsystem), two non-vanishing equations from  (54), and (55) for each subsystem giving the relaxation of the shear stress tensor and the eight non-vanishing coupling equations (2.1). These equations are given in Appendix A. Utilizing these equations, we can solve for the fourteen independent variables, namely the seven variables $\varepsilon$, $\pi_{d}$, $\pi_{od}$, $a_{11}$, $a_{22}$, $a_{33}$ and $a_{23}$ of the first subsystem and the seven corresponding ones for the other. To initialize, we need to specify only the initial values of six variables, namely $\varepsilon$, $\pi_{d}$, and $\pi_{od}$ of the first subsystem and the three corresponding ones for the other subsystem. This is because the variables of the effective metrics are obtained via the algebraic coupling equations.
@@
For numerical simulations in $2+1$-dimensions, we choose parameters that are lead to the first subsystem being strongly self-interacting and the second one being weakly self-interacting. Both subsystems are conformal. We choose the equations of states of the two subsystems so that $\varepsilon/T^{3}=n_{1}=\tilde{\varepsilon}/\tilde{T}^{3}=n_{2}=2$. The shear viscosity and the relaxation times of the subsystems are chosen as shown in Table 1. We proceed to numerically solve the algebraic-differential system (explicitly given in appendix A), such that the total energy density is conserved to a precision of $10^{-6}$.
@@
Figure 3: Evolution of the full system entropy density $\mathcal{S}$ (left) and the subsystem entropy densities $s_{1}$ and $s_{2}$ (right) for fixed energy density $\mathcal{E}_{0}=0.991$ and anisotropy $A1$ given in Table 2. Here we have fixed $\gamma=1$ and $\tilde{\gamma}=-2$. We note that although the full entropy density reaches very close to the global equilibrium value, the subsystem entropy densities are away from the global equilibrium value. The final state is a pseudo-equilibrium state. We can explicitly check that the final state is typically much closer to the global equilibrium state for larger values of $\mathcal{E}_{0}$,
@@ 5.2 Thermalization for large energy densities
  • Choose a value of $\delta:=(a_{11}T-\tilde{a}_{11}\tilde{T})/(a_{11}T+\tilde{a}_{11}\tilde{T})$, where we note that $-1\leq\delta\leq 1$.
@@
Figure 4: Total energy density as a function of $T_{0}$ for typical initial state with $\delta=0$ for the three cases. For the three different anisotropies in Table 1, the curves overlap.
@@
Figure 5: Dependence of the total diagonal anisotropy $\Pi_{d}=\Pi_{22}=-\Pi_{33}$ (left) and off-diagonal anisotropy $\Pi_{od}=\Pi_{23}=\Pi_{32}$ (right) of the system as a function of the energy density of the full system for different initial anisotropy in Table 2.
It can be expected that the entropy of the final pseudo-equilibrium state should have universal characteristics in light of the expectation that the full shear stress tensor governs entropy production and scales with $\mathcal{E}_{0}$ universally in a typical initial state at large $\gamma\mathcal{E}_{0}$. Fig. 6 illustrates the characteristics of the final entropy as a function of the initial anisotropy $\Pi$ (collectively denoting $\Pi_{d}$ and $\Pi_{od}$) of the full system for different choices of the four variables characterizing the initial subsystem shear stress tensors shown in Table 2. We find these features:
@@
Figure 6: Top left: behavior of the final entropy of the full system as a function $\Pi_{d}=\Pi_{22}=-\Pi_{33}$. Top right: how far the initial entropy of the system is from the entropy of the global equilibrium state after anisotropy is turned on. Bottom: entropy production as a function of the total anisotropy of the system. The different curves correspond to the anisotropy in Table 2.
@@ 6 Conclusions
Such a non-linear unitary channel for the perturbative sector with semi-holographic coupling was studied explicitly in Kibe:2023ixa (see also Banerjee:2024ioz ) where a single quantum harmonic oscillator was coupled to a holographic quantum dot described by Jackiw-Teitelboim gravity in the large $N$ limit. In this system, while the entropy of the holographic sector increases, the quantum harmonic oscillator does not lose all of its energy to the black hole. Rather, it goes to a time-dependent state which is determined by the requirement that it decouples from the black hole at late time when the mass of the black hole and the energy of the oscillator become constant. The final time-dependent state also retains a non-isometric copy of the information of the initial state.  [^7] In other semi-holographic scenarios Ecker:2018ucc ; Mondkar:2021qsf , where a single field was coupled to holography, the former lost all its energy to the black hole. Interestingly, when a quantum field is coupled semi-holographically to the black hole, the former retains a residual energy and a non-isometric copy of the initial information. It was shown in Kibe:2023ixa that another copy of the information of the initial state of the quantum oscillator is transferred to the non-linearities of the ringdown process as the black hole mass settles down to a constant value. For a more precise understanding of the quantum statistics of thermalization in large $N$ semi-holography, we need to understand the dynamics of $\mathcal{O}(N^{2})$ quantum oscillators coupled to the black hole via the democratic coupling. It would be interesting to see how the time evolution captures features predicted by the eigenstate thermalization hypothesis via explicit calculations of time-dependent expectation values of operators. We leave this for the future.
@@ Appendix A Explicit expressions
The full determinant discussed around eq. 46 is given by
@@
00 0 $=-1+a_{11}(t)^{2}+\tilde{a}_{11}(t)\sqrt{\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-\tilde{a}_{23}(t)^{2}}\Big{(}\gamma\frac{\tilde{\varepsilon}(t)}{\tilde{a}_{11}(t)^{2}}+\gamma r\Big{(}-\frac{\tilde{\varepsilon}(t)}{\tilde{a}_{11}(t)^{2}}$
@@
00 0 $=1-a_{22}(t)^{2}+\tilde{a}_{11}(t)\sqrt{\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-\tilde{a}_{23}(t)^{2}}\Big{(}\gamma\frac{\left(\tilde{a}_{33}(t)^{2}(\tilde{\varepsilon}(t)-\tilde{\pi}_{d}(t))-2\tilde{a}_{23}(t)\tilde{\pi}_{od}(t)\right)}{2\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-2\tilde{a}_{23}(t)^{2}}$
@@
00 0 $=-a_{23}(t)+\gamma\frac{\tilde{a}_{11}(t)\sqrt{\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-\tilde{a}_{23}(t)^{2}}\left(2\tilde{a}_{22}(t)^{2}\tilde{\pi}_{od}(t)+\tilde{a}_{23}(t)(\tilde{\pi}_{d}(t)-\tilde{\varepsilon}(t))\right)}{2\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-2\tilde{a}_{23}(t)^{2}}$ (66)

### v2

@@ 4 Global equilibrium maximizes the total entropy for the isolated hybrid system
[…]
@@
$\qquad\times a(\lambda)b(\lambda)^{d-1}\tilde{a}(\lambda)\tilde{b}(\lambda)^{d-1}.$
is a constant [^3] Since we are considering homogeneous configurations, the total energy density should be held fixed and the total entropy density should be maximized. It will be clear from the proof that it is sufficient to prove the maximization of entropy in global thermal equilibrium in homogeneous equilibrium as the proof can be implemented locally.. We recall that $a(\lambda)$, $b(\lambda)$, $\tilde{a}(\lambda)$ and $\tilde{b}(\lambda)$ should be obtained by solving the coupling equations sec.˜3.2 (with $\phi=0$) which are fully specified by the subsystem temperatures, $T_{1}(\lambda)$ and $T_{2}(\lambda)$.
@@
For illustration, we can consider the case of two subsystems with conformal equations of states $\epsilon=n_{1}T_{1}^{3}$ and $\tilde{\epsilon}=n_{2}T_{2}^{3}$ in $2+1$-dimensions. We choose $n_{1}=10$ and $n_{2}=1$. In the left panel in Fig.˜2, we note that while one of the subsystem entropies monotonically increases with its effective temperature, that of the other decreases as a consequence of the fixed total energy. With the total energy $\mathcal{E}$ kept constant and setting $\lambda=T_{1}$, the total entropy $\mathcal{S}$ has been plotted alongside $T_{a}-T_{b}$ as a function of $T_{1}$ in the right panel of Fig.˜2. We observe that $T_{a}-T_{b}$ indeed vanishes exactly where $\mathcal{S}$ is maximum. We also note that the total entropy is a concave function and has only one maximum corresponding to the global equilibrium.
@@ 5.1 Hybrid BRSSS setup
$g_{{\mu\nu}}=\begin{pmatrix}-a_{11}(t)^{2}&0&0\\ 0&a_{22}(t)^{2}&a_{23}(t)\\ 0&a_{23}(t)&a_{33}(t)^{2}\end{pmatrix}\hskip 14.22636pt\text{and}\hskip 14.22636pt{\tilde{g}}_{{\mu\nu}}=\begin{pmatrix}-\tilde{a}_{11}(t)^{2}&0&0\\ 0&\tilde{a}_{22}(t)^{2}&\tilde{a}_{23}(t)\\ 0&\tilde{a}_{23}(t)&\tilde{a}_{33}(t)^{2}\end{pmatrix}$. (54)
@@
[…]
@@
In the homogeneous thermal relaxation process in $2+1$-dimensions, the full system is governed by fourteen equations; namely, one Ward identity for each subsystem (the conservation of the energy-momentum tensor gives only one non-vanishing equation for each subsystem), two non-vanishing equations from  (59), and (60) for each subsystem giving the relaxation of the shear stress tensor and the eight non-vanishing coupling equations (2.1). These equations are given in Appendix˜A. Utilizing these equations, we can solve for the fourteen independent variables, namely the seven variables $\varepsilon$, $\pi_{d}$, $\pi_{od}$, $a_{11}$, $a_{22}$, $a_{33}$ and $a_{23}$ of the first subsystem and the seven corresponding ones for the other. To initialize, we need to specify only the initial values of six variables, namely $\varepsilon$, $\pi_{d}$, and $\pi_{od}$ of the first subsystem and the three corresponding ones for the other subsystem. This is because the variables of the effective metrics are obtained via the algebraic coupling equations.
@@
For numerical simulations in $2+1$-dimensions, we choose parameters that are lead to the first subsystem being strongly self-interacting and the second one being weakly self-interacting. Both subsystems are conformal. We choose the equations of states of the two subsystems so that $\varepsilon/T^{3}=n_{1}=\tilde{\varepsilon}/\tilde{T}^{3}=n_{2}=2$. The shear viscosity and the relaxation times of the subsystems are chosen as shown in Table 1. We proceed to numerically solve the algebraic-differential system (explicitly given in appendix˜A), such that the total energy density is conserved to a precision of $10^{-6}$.
@@
Figure 3: Evolution of the full system entropy density $\mathcal{S}$ (left) and the subsystem entropy densities $s_{1}$ and $s_{2}$ (right) for fixed energy density $\mathcal{E}_{0}=0.991$ and anisotropy $A1$ given in Table˜2. The inset plot in the left figure shows the evolution of the subsystem entropy density at early time. Here we have fixed $\gamma=1$ and $\tilde{\gamma}=-2$. We note that although the full entropy density reaches very close to the global equilibrium value, the subsystem entropy densities are away from the global equilibrium value. The final state is a pseudo-equilibrium state. We can explicitly check that the final state is typically much closer to the global equilibrium state for larger values of $\mathcal{E}_{0}$.
@@ 5.2 Thermalization for large energy densities
[…]
@@
Figure 4: Total energy density as a function of $T_{0}$ for typical initial state with $\delta=0$ for the three cases. For the three different anisotropies in Table˜1, the curves overlap.
@@
Figure 5: Dependence of the total diagonal anisotropy $\Pi_{d}=\Pi_{22}=-\Pi_{33}$ (left) and off-diagonal anisotropy $\Pi_{od}=\Pi_{23}=\Pi_{32}$ (right) of the system as a function of the energy density of the full system for different initial anisotropy in Table˜2.
It can be expected that the entropy of the final pseudo-equilibrium state should have universal characteristics in light of the expectation that the full shear stress tensor governs entropy production and scales with $\mathcal{E}_{0}$ universally in a typical initial state at large $\gamma\mathcal{E}_{0}$. Fig.˜6 illustrates the characteristics of the final entropy as a function of the initial anisotropy $\Pi$ (collectively denoting $\Pi_{d}$ and $\Pi_{od}$) of the full system for different choices of the four variables characterizing the initial subsystem shear stress tensors shown in Table˜2. We find these features:
@@
Figure 6: Top left: behavior of the final entropy of the full system as a function $\Pi_{d}=\Pi_{22}=-\Pi_{33}$. Top right: how far the initial entropy of the system is from the entropy of the global equilibrium state after anisotropy is turned on. Bottom: entropy production as a function of the total anisotropy of the system. The different curves correspond to the anisotropy in Table˜2.
@@ 6 Conclusions
Such a non-linear unitary channel for the perturbative sector with semi-holographic coupling was studied explicitly in Kibe:2023ixa (see also Banerjee:2024ioz ) where a single quantum harmonic oscillator was coupled to a holographic quantum dot described by Jackiw-Teitelboim gravity in the large $N$ limit. In this system, while the entropy of the holographic sector increases, the quantum harmonic oscillator does not lose all of its energy to the black hole. Rather, it goes to a time-dependent state which is determined by the requirement that it decouples from the black hole at late time when the mass of the black hole and the energy of the oscillator become constant. The final time-dependent state also retains a non-isometric copy of the information of the initial state.  [^8] In other semi-holographic scenarios Ecker:2018ucc ; Mondkar:2021qsf , where a single field was coupled to holography, the former lost all its energy to the black hole. Interestingly, when a quantum field is coupled semi-holographically to the black hole, the former retains a residual energy and a non-isometric copy of the initial information. It was shown in Kibe:2023ixa that another copy of the information of the initial state of the quantum oscillator is transferred to the non-linearities of the ringdown process as the black hole mass settles down to a constant value.
For a more precise understanding of the quantum statistics of thermalization in large $N$ semi-holography, we need to understand the dynamics of $\mathcal{O}(N^{2})$ quantum oscillators coupled to the black hole via the democratic coupling. In the presence of the effective metric coupling, we expect that only this coupling survives at late time giving an evolution described by the $U[g_{\mu\nu}]\otimes\tilde{U}[\tilde{g}_{\mu\nu}]$ non-linear unitary channel. It would be interesting to see how the time evolution captures features predicted by the eigenstate thermalization hypothesis via explicit calculations of time-dependent expectation values of operators. [^9] For a recent demonstration of how holographic systems realize statistical features of the eigenstate thermalization hypothesis, see Gannouji:2025nmh . We leave this for the future.
@@ Appendix A Explicit expressions
The full determinant discussed around eq.˜51 is given by
@@
$0$ $=-1+a_{11}(t)^{2}+\tilde{a}_{11}(t)\sqrt{\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-\tilde{a}_{23}(t)^{2}}\Big(\gamma\frac{\tilde{\varepsilon}(t)}{\tilde{a}_{11}(t)^{2}}+\gamma r\Big(-\frac{\tilde{\varepsilon}(t)}{\tilde{a}_{11}(t)^{2}}$
@@
$0$ $=1-a_{22}(t)^{2}+\tilde{a}_{11}(t)\sqrt{\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-\tilde{a}_{23}(t)^{2}}\Big(\gamma\frac{\left(\tilde{a}_{33}(t)^{2}(\tilde{\varepsilon}(t)-\tilde{\pi}_{d}(t))-2\tilde{a}_{23}(t)\tilde{\pi}_{od}(t)\right)}{2\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-2\tilde{a}_{23}(t)^{2}}$
@@
$0$ $=-a_{23}(t)+\gamma\frac{\tilde{a}_{11}(t)\sqrt{\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-\tilde{a}_{23}(t)^{2}}\left(2\tilde{a}_{22}(t)^{2}\tilde{\pi}_{od}(t)+\tilde{a}_{23}(t)(\tilde{\pi}_{d}(t)-\tilde{\varepsilon}(t))\right)}{2\tilde{a}_{22}(t)^{2}\tilde{a}_{33}(t)^{2}-2\tilde{a}_{23}(t)^{2}}$ (71)